• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name gilman

    Log entry time 14:59:44 on March 9,2009

    Entry number 265464

    keyword=9 March 2009 day shift summary

    8:00 AM: Ask MCC for raster off.

    8:10 AM: Ask for HARP scan, 100-300 u

    9:10: done raster size check, move to pol 3He / NMR measurement
    JP reports Pol = 63%.

    9:12: request 14 uA beam for pol 3He production.

    9:44 beam goes off as we ask for controlled access for polarity change

    10:20 Hall rad survey done

    10:30 set BB and HRSL magnets to 0.

    -- controlled access rest of shift --

    -- MCC warming up 1L04, expect beam perhaps late swing shift if all goes well.